# Defensive Unlearning with Adversarial Training for Robust Concept Erasure in Diffusion Models

### v2

@@ 7 Conclusion
8 Acknowledgement
Y. Zhang, J. Jia, Y. Zhang, C. Fan, J. Liu, and S. Liu were supported by the National Science Foundation (NSF) Robust Intelligence (RI) Core Program Award IIS-2207052, the NSF CPS Award CNS-2235231, and the Cisco Faculty Research Award.
@@ Appendix A System Prompt for LLM as a Judger
We utilize a Large Language Model (LLM) like GPT-4 to act as a judger to filter out prompts that are relevant to the concept to be unlearned. Since nudity is a global concept (Gandikota et al., 2023a), even if ’nudity’ is not mentioned in the prompt, the generated image conditioned on the prompt might still involve the ’nudity’ concept. As a result, the system prompt for the LLM to act as a prompt judger for the ’nudity’ concept is shown below:
@@ Appendix C Attack Setup
Following the methodology of (Zhang et al., 2023b), we introduce prepended prompt perturbations using tokens, where is designated for nudity unlearning and for both style and object unlearning. To optimize these adversarial perturbations, we sample diffusion time steps and execute UnlearnDiffAtk over iterations with a learning rate of , using the AdamW optimizer. For the classification of generated images, we use NudeNet (Bedapudi, 2019) for nudity unlearning, a ViT-base model (Zhang et al., 2023b) fine-tuned on the WikiArt dataset for style unlearning, and an ImageNet-pretrained ResNet-50 for object unlearning tasks.
@@ Appendix D Other Results and Visualizations for Nudity Unlearning
The metrics used for this evaluation are Attack Success Rate (ASR), Fréchet Inception Distance (FID), and CLIP score. As shown in Tab. A1, although ScissorHands (SH) (Wu and Harandi, 2024) and EraseDiff (ED) (Wu et al., 2024) achieve high unlearning robustness against adversarial prompt attacks, the trade-offs are significant. Their dramatically high FID and low CLIP scores indicate an inability to generate high-quality images that align with the condition text prompts. This is further corroborated by their visualization examples. In Fig. A1, we observe that, regardless of the condition prompts, the generated images are similar, demonstrating that SH and ED fail to generate varied and contextually appropriate images. Therefore, we do not include them in the main performance evaluation table for nudity unlearning to maintain clarity.
@@ Appendix E Other Results and Visualizations for Object Unlearning
In Tab. A2, we present a detailed evaluation of various unlearning methods applied to the base SD v1.4 model for three different objects: Garbage Truck, Parachute, and Tench. The unlearning methods compared include FMN (Forget-Me-Not) (Zhang et al., 2023a), SPM (concept-SemiPermeable Membrane) (Lyu et al., 2023), SalUn (saliency unlearning) (Fan et al., 2023), ED (EraseDiff) (Wu et al., 2024), ESD (erased stable diffusion) (Gandikota et al., 2023a), SH (ScissorHands) (Wu and Harandi, 2024), and our proposed DM unlearning scheme, referred to as AdvUnlearn. As shown, our proposed DM unlearning method, AdvUnlearn, consistently achieves the best unlearning efficacy (around ) with competitive image generation utility. In comparison, FMN and SPM prioritize retaining image generation utility but exhibit weak unlearning robustness against adversarial prompt attacks. Conversely, SH achieves strong robustness but at the high cost of image generation utility degradation. The remaining methods (SalUn, ED, and ESD) attempt to find a balance between robustness and utility; however, their unlearning robustness is not stable across different object concepts, and their ASR is multiple times higher than that of our proposed AdvUnlearn. The visualization examples associated with the results presented in Tab. A2 can be found in Fig. A2. Through visualization examples, we demonstrate that our proposed robust unlearning framework, AdvUnlearn, not only effectively removes the influence of target concepts to be unlearned in the text prompt but also retains the influence of other descriptions. For instance, a garbage truck-unlearned DM equipped with the AdvUnlearn text encoder generates a photo of a parking lot from the text prompt ‘garbage truck in a parking lot.’ Similarly, other object-unlearned DMs with corresponding AdvUnlearn text encoders produce a photo of a desert landscape from the prompt ‘parachute in a desert landscape’ and a photo of a baby in a pond from the prompt ‘baby tench in a pond.’ Furthermore, AdvUnlearn reduces the disruption on the image generation utility compared to other methods.
@@ Appendix F Other Ablation Studies
Number of attack steps. In Fig. A4, we delve into the dynamics of adversarial prompt generation as detailed in (6) within the framework of AdvUnlearn. The graph distinctly illustrates that as we incrementally increase the number of attack steps, there is a notable enhancement in the efficacy of the unlearning process. This improvement is consistently observed up to the threshold of attack steps, where the curve begins to plateau. Interestingly, during this escalation, there is only a marginal decline in image generation utility, suggesting a robust balance between unlearning performance and utility retention. However, surpassing the step mark appears to have a negligible impact on further augmenting unlearning efficacy, indicating an optimal saturation point in the adversarial training process. This insight is crucial for optimizing the efficiency of the AdvUnlearn algorithm, as it delineates a clear boundary for maximizing gains while minimizing resource expenditure.
@@
Robustness gain from AdvUnlearn at different test-time attacks. In Tab. A5, we present the robustness of SD v1.4 post-nudity unlearning when facing different test-time adversarial prompt attacks, UnlearnDiffAtk (Zhang et al., 2023b) (default choice) and P4D (Chin et al., 2023). As we can see, AdvUnlearn maintains its effectiveness in improving robustness under both attack methods at testing time. Notably, the ASR against P4D is even lower than that against UnlearnDiffAtk. This result is expected, as P4D employs the same attack loss (6) used during training for generating adversarial prompts in (9). Therefore, we default to UnlearnDiffAtk for test-time attacks in robustness evaluation due to its unseen nature during training.

### v3

@@ 7 Conclusion
Acknowledgement
Y. Zhang, J. Jia, Y. Zhang, C. Fan, J. Liu, and S. Liu were supported by the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the NSF CAREER Award IIS-2338068, the Cisco Research Award, and the Amazon Research Award for AI in Information Security.
@@ Appendix A System Prompt for LLM as a Judger
We utilize a Large Language Model (LLM) like GPT-4 to act as a judger to filter out prompts that are relevant to the concept to be unlearned. Since nudity is a global concept [19], even if ’nudity’ is not mentioned in the prompt, the generated image conditioned on the prompt might still involve the ’nudity’ concept. As a result, the system prompt for the LLM to act as a prompt judger for the ’nudity’ concept is shown below:
@@ Appendix C Attack Setup
Following the methodology of [23], we introduce prepended prompt perturbations using tokens, where is designated for nudity unlearning and for both style and object unlearning. To optimize these adversarial perturbations, we sample diffusion time steps and execute UnlearnDiffAtk over iterations with a learning rate of , using the AdamW optimizer. For the classification of generated images, we use NudeNet [113] for nudity unlearning, a ViT-base model [23] fine-tuned on the WikiArt dataset for style unlearning, and an ImageNet-pretrained ResNet-50 for object unlearning tasks.
@@ Appendix D Other Results and Visualizations for Nudity Unlearning
The metrics used for this evaluation are Attack Success Rate (ASR), Fréchet Inception Distance (FID), and CLIP score. As shown in Tab. A1, although ScissorHands (SH) [54] and EraseDiff (ED) [53] achieve high unlearning robustness against adversarial prompt attacks, the trade-offs are significant. Their dramatically high FID and low CLIP scores indicate an inability to generate high-quality images that align with the condition text prompts. This is further corroborated by their visualization examples. In Fig. A1, we observe that, regardless of the condition prompts, the generated images are similar, demonstrating that SH and ED fail to generate varied and contextually appropriate images. Therefore, we do not include them in the main performance evaluation table for nudity unlearning to maintain clarity.
@@ Appendix E Other Results and Visualizations for Object Unlearning
In Tab. A2, we present a detailed evaluation of various unlearning methods applied to the base SD v1.4 model for three different objects: Garbage Truck, Parachute, and Tench. The unlearning methods compared include FMN (Forget-Me-Not) [20], SPM (concept-SemiPermeable Membrane) [28], SalUn (saliency unlearning) [27], ED (EraseDiff) [53], ESD (erased stable diffusion) [19], SH (ScissorHands) [54], and our proposed DM unlearning scheme, referred to as AdvUnlearn. As shown, our proposed DM unlearning method, AdvUnlearn, consistently achieves the best unlearning efficacy (around ) with competitive image generation utility. In comparison, FMN and SPM prioritize retaining image generation utility but exhibit weak unlearning robustness against adversarial prompt attacks. Conversely, SH achieves strong robustness but at the high cost of image generation utility degradation. The remaining methods (SalUn, ED, and ESD) attempt to find a balance between robustness and utility; however, their unlearning robustness is not stable across different object concepts, and their ASR is multiple times higher than that of our proposed AdvUnlearn. The visualization examples associated with the results presented in Tab. A2 can be found in Fig. A2. Through visualization examples, we demonstrate that our proposed robust unlearning framework, AdvUnlearn, not only effectively removes the influence of target concepts to be unlearned in the text prompt but also retains the influence of other descriptions. For instance, a garbage truck-unlearned DM equipped with the AdvUnlearn text encoder generates a photo of a parking lot from the text prompt ‘garbage truck in a parking lot.’ Similarly, other object-unlearned DMs with corresponding AdvUnlearn text encoders produce a photo of a desert landscape from the prompt ‘parachute in a desert landscape’ and a photo of a baby in a pond from the prompt ‘baby tench in a pond.’ Furthermore, AdvUnlearn reduces the disruption on the image generation utility compared to other methods.
@@ Appendix F Other Ablation Studies
Robustness gain from AdvUnlearn at different test-time attacks. In Tab. A5, we present the robustness of SD v1.4 post-nudity unlearning when facing different test-time adversarial prompt attacks, UnlearnDiffAtk [23] (default choice) and P4D [24]. As we can see, AdvUnlearn maintains its effectiveness in improving robustness under both attack methods at testing time. Notably, the ASR against P4D is even lower than that against UnlearnDiffAtk. This result is expected, as P4D employs the same attack loss (6) used during training for generating adversarial prompts in (9). Therefore, we default to UnlearnDiffAtk for test-time attacks in robustness evaluation due to its unseen nature during training.